Apply the method of variation of parameters to solve the following differential equation
d^2y/dx^2+ x^2/y=secnx , where n is a constant
1
Expert's answer
2020-12-15T19:43:53-0500

"\displaystyle \frac{\mathrm{d}^2y}{\mathrm{d}x^2} + n^2y = \sec(nx)

\\ \textsf{The solution to the above equation is}
\\ y = y_c + y_p \\
\textsf{where}\, y_c \, \textsf{is the complementary factor and}
\\ y_p\, \textsf{ is the particular integral}
\\ \textsf{The auxiliary equation is} \,m^2 + n^2 = 0 \\
 
\begin{aligned}
 (m - jn)(m + jn) &= 0 \\ 
\therefore m &= \pm jn \hspace{1cm} \{\textsf{where}\, j\, \textsf{is a complex number}\}
\end{aligned} \\

\textsf{Recall that if the solution of the auxiliary equation}\\
\textsf{of a second-order differential equation is}\\
m = \alpha \pm j\beta, \textsf{the general solution is} \\

y = e^{\alpha x}(C_1\cos{\beta x} + C_2\sin{\beta x})\\

\therefore y_c = C_1\cos(nx) + C_2\sin(nx)\\

\textsf{The Wronskian of the two solutions is} \\
W(x) = \begin{vmatrix} 
\cos(nx) & \sin(nx)\\ 
\frac{\mathrm{d}}{\mathrm{d}x}(\cos(nx)) & \frac{\mathrm{d}}{\mathrm{d}x}(\sin(nx)) 
\end{vmatrix} = 

\begin{vmatrix} 
\cos(nx) & \sin(nx)\\
 -n\sin(nx) & n\cos(nx)
 \end{vmatrix} \\ 

\begin{aligned}
&= n\cos^2(nx) + n\sin^2(nx)\\
&= n(\cos^2(nx) + \sin^2(nx)) \\
&= n
\end{aligned}\\

\therefore\textsf{ Our particular solution will be given by}\\ 
y_p = V_1(x)\cos(nx) + V_2(x)\sin(nx)\\

\textsf{Where}\, V_1(x) = -\int \frac{r(x)\sin(nx)}{W(x)} \, \mathrm{d}x,\, V_2(x) = \int \frac{r(x)\cos(nx)}{W(x)}\, \mathrm{d}x \, \textsf{and} \, \\

\begin{aligned} 
V_1(x) &= -\int \frac{r(x)\sin(nx)}{W(x)} \, \mathrm{d}x \\
&= -\int \frac{\sec(nx)\sin(nx)}{n}\, \mathrm{d}x\\
&= -\int \frac{\tan(nx)}{n}\, \mathrm{d}x\\
&= -\frac{\ln(\sec(nx))}{n^2} + C
\end{aligned} \\

\begin{aligned} 
V_2(x) &= \int \frac{r(x)\cos(nx)}{W(x)}\, \mathrm{d}x \\
&= \int \frac{\sec(nx)\cos(nx)}{n}\, \mathrm{d}x\\
&= \int \frac{1}{n}\, \mathrm{d}x\\
&= \frac{x}{n} + C
\end{aligned} \\
\textsf{The constant terms of the integration can be}
\\\textsf{ignored since we are trying to find a non-constant}
\\\textsf{solution to the differential equation}

\\ \begin{aligned}
\therefore y &= y_c + y_p \\
&=C_1\cos(nx) + C_2\sin(nx) - \frac{1}{n^2}\ln(\sec(nx))\cos(nx) +\frac{x}{n}\sin(nx)
\end{aligned}"